OpenStudy (anonymous):

Click an item in the list or group of pictures at the bottom of the problem and holding your mouse button down, drag it into the correct position in the area. Release your mouse button when the item is place. If you change your mind, double-click one item to remove it, or click the Start Over button to clear all of your answers. Write the equation for the following relation. D = {(x, y): (3, 9), (5, 25), (-2, 4), . . .}

OpenStudy (helder_edwin):

(3,9) 9=3^2 (5,25) 25=5^2 (-2,4) 4=(-2)^2

OpenStudy (anonymous):

Thank you...you made it understandable...the text take me around the world!!!

OpenStudy (helder_edwin):

u r welcome